Round to the nearest hundredth.
The two legs of a right triangle measure 6 in. and 2 in. Find the length of the hypotenuse.

Found 2 solutions by Alan3354, ewatrrr:
Answer by Alan3354(69443)
Answer by ewatrrr(24785)


Hi

The two legs of a right triangle measure 6 in. and 2 in.
Find the length of the hypotenuse.

Pythagorean Theorem:  
A^2 + B^2 = C^2
  6^2 + 2^2 = 40 
      c^2 = 40     Tossing out the negative solution for length
      c = 6.32 in  Rounded to the nearest hundredth.

Important You feel comfortable using Your calculator
Wish You the Best in your Studies.